OpenStudy (anonymous):

a particle is constrained to move along a straight line through O. it starts initially at D, which is a fixed distance from O in the positive direction&moves at a constant velocity of 4ms^-1 for 15 sec. until; it is 100m from O for a furhter 25 sec, after which it travels in the opposite direction for 20 sec. the total distance is 180m. How far from O does the particle stop? how do i draw a velocity–time graph? and what is the final displacement of the particle from its initial position?

OpenStudy (anonymous):

Refer to picture.

OpenStudy (anonymous):

To start with, it travels at 4 m/s for 15 seconds, so how far does it travel in that time? Distance = velocity * time, so it travels 4*15 = 60 m.

OpenStudy (anonymous):

It travels 60 meters and ends up 100 meters from O, which means that point D is 100-60 = 40 meters from O.

OpenStudy (anonymous):

The next thing to figure out is how fast it is traveling when it goes backwards. Well, we know that in total it travels 180 meters, that's given. It initially traveled 60 meters, so that means that it travels backwards for 180 - 60 = 120 meters.

OpenStudy (anonymous):

Velocity = distance/time, so it is traveling at a speed of 120/20 = 6 m/s

OpenStudy (anonymous):

That's a picture of the point traveling on the line.

OpenStudy (anonymous):

The graph will look a lot different. It needs to have an x and y axis.

OpenStudy (anonymous):

im sorry but im confused dunno what you did after 100-60=40 then im lost :(

OpenStudy (anonymous):

is it constant, so do i draw a straight line?